# Mapping the Cosmic Gamma-ray Horizon: The 1CGH Catalogue of Fermi-LAT detections above 10 GeV

### v2

@@ Appendix B Long Tables: Best Candidates for optical Observations
In this section, we present tables highlighting the best candidates for optical follow-up observations from the 1CGH catalogue. These sources have been selected based on their potential to provide insights about the CGH. Two categories of sources are prioritized: Table 31 those without redshift information but with clear optical, infrared, or radio associations, and Table 42 those with uncertain or lower-limit redshifts where significant $\gamma$-ray absorption is predicted. These candidates will refine the redshift completeness of the 1CGH catalogue. Each source selection is based -respectively- on its highest-energy bin ($\rm E^{bin}_{max}$), and on its calculated optical depth $\rm\tau_{(Ebin)}$, as described in sec. 3.4. Table 3 provides the IR counterparts from the AllWISE catalogue (Cutri et al., 2021) which are meant to help guide the optical observations [^12] For 4FGLJ0351.2-6103, the AllWISE IR counterpart is tentative; and for 4FGLJ2127.6-5959, a tentative IR counterpart is close by NGC7059..
[…]
Name R.A.(deg) Dec.(deg) 4FGLdr4 ASSOC1-4FGL TS $\rm E^{bin}_{max}$ Rmag AllWISE
@@
Table 4: This table lists the sources in the catalogue with uncertain redshift information, ordered by their $\tau_{(E,z)}$ values calculated at $\rm E^{bin}_{max}$. The first columns show the source names, R.A., and Dec. in degrees (J2000), respectively. The fourth column presents redshift information from the literature, with a ‘(?)’ flag indicating uncertain or photometric values and a ‘>’ symbol denoting lower limits. The ‘z-origin’ column specifies the literature reference for the redshift, using the short names described in Section 3.1. The ‘b(deg)’ column shows the galactic latitude in degrees, followed by the associated ‘4FGL-DR4’ name. The ‘$\rm TS^{1CGH}_{10GeV}$’ column provides the Test Statistic value derived for 1CGH sources detected above 10 GeV, integrating over 16 years of Fermi-LAT observations. The columns ‘$\rm E_{max}^{bin}$’ and ‘$\rm\tau_{Ebin}$’ list the largest energy bin detected and the corresponding EBL optical depth $\tau_{(E,z)}$, respectively. For sources whose Name’ begins with 4FGLJ, the R.A. and Dec. correspond to the associated counterpart position(see the ‘ASSOC1-4FGL’ column in, Ballet et al., 2023).
[…]

### v3

@@ Appendix B Long Tables: Best Candidates for optical Observations
In this section, we present tables highlighting the best candidates for optical follow-up observations from the 1CGH catalogue. These sources have been selected based on their potential to provide insights about the CGH. Two categories of sources are prioritized: Table 31 those without redshift information but with clear optical, infrared, or radio associations, and Table 42 those with uncertain or lower-limit redshifts where significant $\gamma$-ray absorption is predicted. These candidates will refine the redshift completeness of the 1CGH catalogue. Each source selection is based -respectively- on its highest-energy bin ($\rm E^{bin}_{max}$), and on its calculated optical depth $\rm\tau_{(Ebin)}$, as described in sec. 3.4. Table 3 provides the IR counterparts from the ‘Wide-field Infrared Survey Explorer’ AllWISE catalogue (Cutri et al., 2013, 2021) which are meant to help guide optical observations [^12] For 4FGLJ0351.2-6103, the AllWISE IR counterpart is tentative; and for 4FGLJ2127.6-5959, a tentative IR counterpart is close by NGC7059..
Table 3: This table lists the sources in the catalogue without redshift information, ordered by their $\rm E^{bin}_{max}$, and where we could retrieve clear IR/Optical counterparts for observation. The first three columns show the source names, right ascension (R.A.), and declination (Dec.) in degrees (J2000), respectively. The column ‘4FGL-DR4’ and ‘ASSOC1-4FGL’ shows the source and counterpart names reported in 4FGL-DR4 catalogue. Next, the column ‘TS’ provides the Test Statistic value derived for 1CGH sources detected above 10 GeV, integrating over 16 years of Fermi-LAT observations. The column ‘$\rm E_{max}^{bin}$’ list the highest energy bin detected. The column ‘Rmag’ provides the R band magnitude retrieved using SSDC Sky Explorer.(https://tools.ssdc.asi.it/). The last column, ‘AllWISE’ provides the IR counterpart from the AllWISE catalogue (cases marked with a ‘*’ sign represent a tentative association and should be considered with care). For sources whose ’Name’ begins with 4FGLJ, the R.A. and Dec. correspond to the counterpart position (‘ASSOC1’, Ballet et al., 2023).
Name R.A.(deg) Dec.(deg) 4FGL-DR4 ASSOC1-4FGL TS $\rm E^{bin}_{max}$ Rmag AllWISE
@@
Table 4: This table lists the sources in the catalogue with uncertain redshift information, ordered by their $\tau_{(E,z)}$ values calculated at $\rm E^{bin}_{max}$. The first columns show the source names, R.A., and Dec. in degrees (J2000), respectively. The fourth column presents redshift information from the literature, with a ‘(?)’ flag indicating uncertain or photometric values and a ‘>’ symbol denoting lower limits. The ‘z-origin’ column specifies the literature reference for the redshift, using the short names described in Section 3.1. The ‘b(deg)’ column shows the galactic latitude in degrees, followed by the associated ‘4FGL-DR4’ name. The ‘TS’ column provides the Test Statistic value derived for 1CGH sources detected above 10 GeV, integrating over 16 years of Fermi-LAT observations. The columns ‘$\rm E_{max}^{bin}$’ and ‘$\rm\tau_{Ebin}$’ list the largest energy bin detected and the corresponding EBL optical depth $\tau_{(E,z)}$, respectively. For sources whose Name’ begins with 4FGLJ, the R.A. and Dec. correspond to the associated counterpart position(see the ‘ASSOC1-4FGL’ column in, Ballet et al., 2023).
Name R.A. (deg) Dec. (deg) z z-origin b(deg) 4FGL-DR4 TS $\rm E^{bin}_{max}$ $\rm\tau_{Ebin}$